Find the value of y so that the line passing through the two points has a given slope
(2,y)(11,8)

Answer by nerdybill(7384)

Find the value of y so that the line passing through the two points has a given slope
(2,y)(11,8)
.
You didn't state what the "given slope" was...
But, let's call it m
then
m = (y2 - y1)/(x2 - x1)
m = (8 - y)/(11 - 2)
m(11 - 2) = (8 - y)
9m = 8 - y
-9m = -8 + y
8-9m = y
.
So, plug in the "given slope" (m) above, and you'll find y